OpenStudy (anonymous):

What is the d value in the following arithmetic sequence? ... -7, -2, 3, 8, … +5

OpenStudy (anonymous):

+5 was my answer

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay :) thanks

OpenStudy (anonymous):

This is an arithmetic sequence given in the form: a, a + d, a + 2d...and so on. To find d, we can subtract a + d (the second term) with a (the first term). (a + d) - (a) = d = -2 - (-7) = -2 + 7 = 5

OpenStudy (anonymous):

@anthony~